1. The rate constant for a certain reaction is k = 5.70×10−3 s−1 . If the initial reactant concentration was 0.750 M, what will the concentration be after 13.0 minutes?

2. A zero-order reaction has a constant rate of 4.40×10−4 M/s. If after 80.0 seconds the concentration has dropped to 1.00×10−2 M, what was the initial concentration?
Answer #1

From the units of rate constant s-1 it can be observed that it is a first order reaction

ln(0.750/b)=5.70*10-3s-1*13*60 sec (one min=60 sec)

ln(0.750/b)=4.446

0.750/b=e4.446

0.750/b=85.28512

b=0.750M/85.28512

b=0.008794 M